# PASCAL: Precise and Efficient ANN- SNN Conversion using Spike Accumulation and Adaptive Layerwise Activation

### v1

@@ Abstract
Spiking Neural Networks (SNNs) have been put forward as an energy-efficient alternative to Artificial Neural Networks (ANNs) since they perform sparse Accumulate operations instead of the power-hungry Multiply-and-Accumulate operations. ANN-SNN conversion is a widely used method to realize deep SNNs with accuracy comparable to that of ANNs. Bu et al. recently proposed the Quantization-Clip-Floor-Shift (QCFS) activation as an alternative to ReLU to minimize the accuracy loss during ANN-SNN conversion. Nevertheless, SNN inferencing requires a large number of timesteps to match the accuracy of the source ANN for real-world datasets. In this work, we propose PASCAL, which performs ANN-SNN conversion in such a way that the resulting SNN is mathematically equivalent to an ANN with QCFS-activation, thereby yielding similar accuracy as the source ANN with minimal inference timesteps. In addition, we propose a systematic method to configure the quantization step of QCFS activation in a layerwise manner, which effectively determines the optimal number of timesteps per layer for the converted SNN. Our results show that the ResNet-34 SNN obtained using PASCAL achieves an accuracy of $\approx$74on ImageNet with a 64$\times$ reduction in the number of inference timesteps compared to existing approaches.
Machine Learning, ICML
@@ 1 Introduction
Deep Artificial Neural Networks (ANNs) have revolutionized the field of machine learning by achieving unprecedented success for various computer vision and natural language processing tasks (LeCun et al., 2015). Despite their superhuman capabilities, there exists a significant computational efficiency divide between ANNs and the human brain. In recent years, Spiking Neural Networks (SNNs) have emerged as a plausible energy-efficient alternative to ANNs (Roy et al., 2019; Tavanaei et al., 2019; Guo et al., 2023). The fundamental building block of an SNN, namely, the spiking neuron, encodes input information using binary spiking events over time. The sparse spike-based computation and communication capability of SNNs have been exploited to obtain improved energy efficiency in neuromorphic hardware implementations (Davies et al., 2021).
@@
Training methodologies for SNNs can be classified primarily into two approaches, namely, direct training using spike-based algorithms and ANN-SNN conversion. Direct SNN training using spike-based error backpropagation (Lee et al., 2016; Wu et al., 2018; Shrestha & Orchard, 2018; Neftci et al., 2019; Lee et al., 2020) has been shown to yield competitive accuracy on complex vision datasets such as ImageNet (Rathi et al., 2020; Deng et al., 2022). It is important to note that SNNs require multiple forward passes, as inputs are fed sequentially over a certain number of timesteps. As a result, performing backpropagation on SNNs is both compute- and memory-intensive because of the need to accumulate error gradients over time. Despite being computationally prohibitive, backpropagation methods have been shown to yield competitive accuracy, with inference latencies typically less than a few tens of timesteps (Deng et al., 2022).
@@
ANN-SNN conversion, on the other hand, eliminates the need for direct SNN training by using the pre-trained ANN weights and replacing the nonlinear activation (for instance, ReLU) with the equivalent spiking non-linearity (integrate-and-fire). The ANN activation is effectively mapped to the average firing rate of the corresponding spiking neuron. The conversion methods have been shown to provide accuracy comparable to that of the original ANN, albeit with longer inference latencies typically spanning hundreds to thousands of timesteps (Sengupta et al., 2019). This is because the correlation between the ANN activation and spiking neuronal firing rate improves with increasing number of timesteps during SNN inference (Cao et al., 2015). Several prior works have proposed techniques to optimize the number of inference timesteps (Diehl et al., 2015; Rueckauer et al., 2017; Sengupta et al., 2018, 2019; Han & Roy, 2020; Deng & Gu, 2021). Most notably, Bu et al. proposed a novel activation function, namely, the Quantization-Clip-Floor-Shift (QCFS) activation, as a replacement for the ReLU activation during ANN training. The QCFS function, with pre-determined quantization step $L$, was formally shown to minimize the expected ANN-SNN conversion error. We note that the QCFS activation works well in practice for CIFAR-10, but for CIFAR-100 and ImageNet, the ANN accuracy can only be matched when the corresponding SNN is inferred for $\sim$1024 timesteps despite being trained using only $4$ or $8$ quantization steps (Bu et al., 2023). In effect, the number of SNN inference timesteps needed to match the ANN accuracy depends on the complexity of the dataset.
@@
Our work addresses this issue using a precise and mathematically grounded ANN-SNN conversion method, in which the number of inference timesteps depends only on the structure of the source ANN, and not on the dataset. We utilize the integrate-and-fire spiking neuron with soft reset (Rueckauer et al., 2017; Han et al., 2020), a technique that is critical for achieving near-zero conversion loss. Next, we address a key limitation of SNNs, which is their inability to handle negative inputs effectively, since spikes are commonly treated as binary values representing either 0 or 1. We demonstrate that inhibitory (or negative) spikes are essential for ensuring the mathematical correctness of our method. Finally, we propose an algorithm to determine the optimal quantization step for QCFS activation per layer by computing a statistical layer-sensitivity metric. This enables an optimal trade-off between computational efficiency and accuracy. In summary, we propose PASCAL, an accurate and efficient ANN-SNN conversion methodology, which has the following twofold contributions.
@@
  • Adaptive Layerwise QCFS Activation. We propose a statistical metric to systematically analyze and categorize layers of a DNN to selectively use higher precision (or QCFS steps) for certain layers, and lower precision for others. The DNN layers with lower precision need to be unrolled for fewer timesteps, thereby, improving the computational efficiency during SNN inference. This is in contrast to early-exit SNNs, in which all the layers are executed for the same number of timesteps, which in turn depends on the input complexity (Srinivasan & Roy, 2021; Li et al., 2023b).
@@ 2.1 ANN-SNN Conversion
ANN-SNN conversion has been shown to be a promising approach for building deep SNNs yielding high enough accuracy for complex visual image recognition (Cao et al., 2015; Diehl et al., 2015; Pérez-Carrasco et al., 2013; Rueckauer et al., 2016, 2017; Sengupta et al., 2019) and natural language processing tasks (Diehl et al., 2016). The conversion algorithms incur the following steps.
@@
Efficient ANN-SNN conversion requires careful threshold initialization at every layer of the network so that the spiking rates are proportional to the ReLU activations. Diehl et al. proposed a model-based threshold balancing scheme, wherein the firing thresholds are determined solely based on ANN activations. Sengupta et al. proposed an improved data-based conversion scheme that additionally uses SNN spiking statistics for threshold initialization. However, the aforementioned methods were susceptible to information loss during SNN inference due to the use of hard reset spiking neurons, wherein the membrane potential is reset to zero in the event of a spike. Han et al. demonstrated near lossless low-latency ANN-SNN conversion with soft reset neurons, wherein the potential is decreased by an amount equal to the neuronal spiking threshold. Subsequently, Bu et al. showed that the threshold voltage could be learned, instead of configuring it heuristically post training, by replacing ReLU with the Quantization-Clip-Floor-Shift (QCFS) activation, which further minimizes the conversion error.
2.2 Quantization-Clip-Floor-Shift (QCFS) Activation
The QCFS function, $\widehat{h}$, as proposed by Bu et al., can be formulated as
where $z^{l}$ is the input to the activation layer, $L$ is the quantization step, and $\lambda^{l}$ is the trainable parameter for the activation threshold. At the end of training, $\lambda^{l}=\theta_{n}$, where $\theta_{n}$ is the final threshold that is subsequently used for SNN inference. Bu et al. showed that the expectation of conversion error reaches $\bm{0}$ when the shift amount, $\bm{\varphi}$, is $\bm{\frac{1}{2}}$.
@@ 2.3 Temporally Efficient SNN Inference
In parallel with improving the ANN-SNN conversion process, there have been efforts to speed up SNN inference by optimizing the number of timesteps needed to achieve the best accuracy. Early-exit inference methods, which incorporate auxiliary classifiers at various stages in the model, have been shown to improve the inference efficiency of both ANNs and SNNs (Panda et al., 2016; Huang et al., 2018; Srinivasan & Roy, 2021)Li et al. proposed a unique early-exit approach for SNNs, wherein the number of inference timesteps is modulated based on the input complexity. In this work, we propose a novel methodology to achieve both precise and temporally efficient ANN-SNN conversion.
@@ 3.1 Precise ANN-SNN Conversion (PASC)
Integrate-and-Fire (IF) layer. We consider the general case where the quantization step $L_{n-1}$ of the previous layer is different from that of the current layer ($L_{n}$). We initialize the layerwise threshold $\theta^{*}=\frac{\theta_{n}}{L_{n}}$, where $\theta_{n}$ is obtained from ANN training (refer to Section 2.2). The initial membrane potential is set to $\frac{\theta^{*}}{2}$, similar to prior works (Bu et al., 2023). We outline the procedure in Algorithm 1 for a generic layer, and Algorithm 2 for the input layer. We formally prove in Section 3.2 that the proposed algorithm enables precise ANN-SNN conversion. The algorithm for a generic layer is described below.
@@
    For the first $L_{n-1}$ timesteps, we perform the IF operation (detailed in 2.1) using the spike train obtained from the preceding layer. We use a spike counter to record the number of spikes produced at this stage.
@@
The output spike train generated by the PASC algorithm is effectively unrolled across $L_{n}$ timesteps by the subsequent layer. Therefore, we preserve the fact that after each layer, the number of inference timesteps is equal to the value of the quantization step, i.e. $T=L_{n}$. We note that Algorithm 2 is similar to, albeit simpler than, Algorithm 1, except that it receives the input image containing real-valued pixels (rather than binary spikes) over time. Feeding the input image to the SNN, without encoding the pixel intensities as spike trains, has been adopted in prior works to minimize accuracy loss during inference (Li et al., 2023b; Bu et al., 2023).
@@
9:     end if
10:  end for
11:  Set $mem(0)$ as the final membrane potential after the previous loop.
12:  for $t\in\{1,2,\dots max(L_{n-1},L_{n})-1\}$ do
13:     if $mem(t)\geq\theta^{*}$ then
14:        $s=s+\frac{\theta_{n}}{L_{n}}$ // Increment on Excitatory spike
15:        $mem(t)=mem(t)-\theta^{*}$
16:     else if $mem(t)<0$ then
17:        $s=s-\frac{\theta_{n}}{L_{n}}$ // Decrement on Inhibitory spike
18:        $mem(t)=mem(t)+\theta^{*}$
19:     end if
20:  end for
21:  Initialize a new SNN with threshold $\theta^{*}=\frac{\theta_{n}}{L_{n}}$, and initial membrane potential $mem(0)=s$, where $s$ is the spike count (scaled by $\theta^{*}$) computed previously.
22:  Initialize an empty tensor $s^{l}$, with dimension $L_{n}\times N\times H\times W$, to generate the final spike train.
23:  for $t\in\{1,2,\dots L_{n}\}$ do
24:     $s^{l}(t)=H(mem(t-1)-\theta^{*})\cdot\theta^{*}$, where $s^{l}(t)$ is the spike output scaled by $\theta^{*}$, and $H$ is the Heaviside step function.
25:     $mem(t)=mem(t-1)-s^{l}(t)$
26:  end for
27:  return $s^{l}$, which will be the input for the subsequent MatMul layer.
@@
6:     $s^{l}(t)=H(mem(t-1)-\theta^{*})\cdot\theta^{*}$, where $s^{l}(t)$ is the spike output scaled by $\theta^{*}$, and $H(x)$ is the Heaviside step function.
7:     $mem(t)=mem(t-1)-s^{l}(t)$
8:  end for
9:  return $s^{l}$
MatMul layers. The MatMul layer performs either a convolution or a fully-connected matrix multiplication, followed by batch normalization. These operations are generally performed using Multiply-and-Accumulate (MAC) operations during ANN inference. It is important to note that the MatMul layers receive spike trains over $L_{n}$ timesteps from the preceding IF layer during SNN inference. The binary spike at each timestep is represented by $\{0,1\}\cdot\theta^{*}$, where the threshold $\theta^{*}=\frac{\theta_{n}}{L_{n}}$. We would like to point out that the quantity obtained by scaling a spike of unit magnitude with the threshold is referred to as postsynaptic potential in the literature (Bu et al., 2023). Therefore, all computations in MatMul layers are unrolled over time and performed on binary input. The former MAC operations are replaced by repeated additions carried out on the spike train.
@@ 3.3 AL: Adaptive Layerwise QCFS Activation
3:  Compute the Skewness $g$ (refer Equation 4), the Van Der Eijk’s Agreement $A$ (refer Section 3.3.1) , and the kurtosis $K$ (refer Equation 5), using the histogram $\mathbf{H}$.
@@ 3.3.1 Van Der Eijk’s Agreement
We use Van Der Eijk’s agreement (Eijk, 2001) to determine how many bins in the histogram $\mathbf{H}$ have a sufficient frequency. It is specified by
@@ 3.3.2 Skewness
Skewness is defined as the deviation of the given distribution from a normal distribution (Joanes & Gill, 1998). It is given by
@@
We use the metric $g^{2}+1$, where $g$ is as defined in Equation 4, as adopted in prior works (Tarbă et al., 2022).
@@ 3.3.3 Kurtosis
where $k_{4}$ is the fourth moment, $k_{2}$ is the second moment, $n$ is the number of samples and $x_{i}$ are the samples whose mean is $\bar{x}$  (Joanes & Gill, 1998). We interpret kurtosis in the following manner.
@@ 3.3.5 Adaptive Layerwise QCFS Activation based Training Methodology
We estimate the AL metric for all the QCFS layers of a pretrained ANN. We use 1-dimensional clustering (Grønlund et al., 2018) to divide the layers into $\chi$ clusters based on the AL metric values, where $\chi$ is a user-defined parameter. The goal is to group QCFS layers with similar AL metric values and assign a suitable quantization step. We subsequently retrain the source ANN from the beginning. This approach is inspired by the Lottery Ticket Hypothesis  (Frankle & Carbin, 2019); the difference being that our subnetwork is not obtained by pruning, but rather by activation quantization. We employ a higher quantization step for the initial epochs to maximize the effect of the learning rate scheduler  (Loshchilov & Hutter, 2016). We then configure the quantization step per layer and incrementally fine-tune the model until training converges, as outlined in Algorithm 4.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Deep Artificial Neural Networks (ANNs) have revolutionized the field of machine learning by achieving unprecedented success for various computer vision and natural language processing tasks (lecun2015deep). Despite their superhuman capabilities, there exists a significant computational efficiency divide between ANNs and the human brain. In recent years, Spiking Neural Networks (SNNs) have emerged as a plausible energy-efficient alternative to ANNs (roy2019towards; tavanaei2019deep; guo2023direct). The fundamental building block of an SNN, namely, the spiking neuron, encodes input information using binary spiking events over time. The sparse spike-based computation and communication capability of SNNs have been exploited to obtain improved energy efficiency in neuromorphic hardware implementations (davies2021advancing).
@@
Training methodologies for SNNs can be classified primarily into two approaches, namely, direct training using spike-based algorithms and ANN-SNN conversion. The efficiency of these methods is compared in A.11. Direct SNN training using spike-based error backpropagation (lee2016training; wu2018spatio; shrestha2018slayer; neftci2019surrogate; lee2020enabling) has been shown to yield competitive accuracy on complex vision datasets such as ImageNet (rathi2020enabling; deng2022temporal). It is important to note that SNNs require multiple forward passes, as inputs are fed sequentially over a certain number of timesteps. As a result, performing backpropagation on SNNs is both compute- and memory-intensive because of the need to accumulate error gradients over time. Despite being computationally prohibitive, backpropagation methods have been shown to yield competitive accuracy, with inference latencies typically less than a few tens of timesteps (deng2022temporal).
@@
ANN-SNN conversion, on the other hand, eliminates the need for direct SNN training by using the pre-trained ANN weights and replacing the nonlinear activation (for instance, ReLU) with the equivalent spiking non-linearity (integrate-and-fire). The ANN activation is effectively mapped to the average firing rate of the corresponding spiking neuron. The conversion methods have been shown to provide accuracy comparable to that of the original ANN, albeit with longer inference latencies typically spanning hundreds to thousands of timesteps (sengupta2019going). This is because the correlation between the ANN activation and spiking neuronal firing rate improves with increasing number of timesteps during SNN inference (cao2015spiking). Several prior works have proposed techniques to optimize the number of inference timesteps (diehl2015fast; rueckauer2017conversion; Sengupta2018Going; sengupta2019going; han2020deep; deng2021optimal). Most notably, bu2023optimal proposed a novel activation function, namely, the Quantization-Clip-Floor-Shift (QCFS) activation, as a replacement for the ReLU activation during ANN training. The QCFS function, with pre-determined quantization step $L$, was formally shown to minimize the expected ANN-SNN conversion error. We note that the QCFS activation works well in practice for CIFAR-10, but for CIFAR-100 and ImageNet, the ANN accuracy can only be matched when the corresponding SNN is inferred for $\sim$1024 timesteps despite being trained using only $4$ or $8$ quantization steps (bu2023optimal). In effect, the number of SNN inference timesteps needed to match the ANN accuracy depends on the complexity of the dataset.
@@
Our work addresses this issue using a precise and mathematically grounded ANN-SNN conversion method, in which the number of inference timesteps depends only on the structure of the source ANN, and not on the dataset. We utilize the integrate-and-fire spiking neuron with soft reset (rueckauer2017conversion; han2020rmp), a technique that is critical for achieving near-zero conversion loss. Next, we address a key limitation of SNNs, which is their inability to handle negative inputs effectively, since spikes are commonly treated as binary values representing either 0 or 1. We demonstrate that inhibitory (or negative) spikes are essential for ensuring the mathematical correctness of our method. Finally, we propose an algorithm to determine the optimal quantization step for QCFS activation per layer by computing a statistical layer-sensitivity metric. This enables an optimal trade-off between computational efficiency and accuracy. In summary, we propose PASCAL, an accurate and efficient ANN-SNN conversion methodology, which has the following twofold contributions.
@@
  • Adaptive Layerwise QCFS Activation. We propose a statistical metric to systematically analyze and categorize layers of a DNN to selectively use higher precision (or QCFS steps) for certain layers, and lower precision for others. The DNN layers with lower precision need to be unrolled for fewer timesteps, thereby, improving the computational efficiency during SNN inference. This is in contrast to early-exit SNNs, in which all the layers are executed for the same number of timesteps, which in turn depends on the input complexity (srinivasan2021bloctrain; li2023seenn).
@@ 2.1 ANN-SNN Conversion
ANN-SNN conversion has been shown to be a promising approach for building deep SNNs yielding high enough accuracy for complex visual image recognition (cao2015spiking; diehl2015fast; perez2013mapping; rueckauer2016theory; rueckauer2017conversion; sengupta2019going) and natural language processing tasks (diehl2016conversion). The conversion algorithms incur the following steps.
@@
Efficient ANN-SNN conversion requires careful threshold initialization at every layer of the network so that the spiking rates are proportional to the ReLU activations. diehl2015fast proposed a model-based threshold balancing scheme, wherein the firing thresholds are determined solely based on ANN activations. sengupta2019going proposed an improved data-based conversion scheme that additionally uses SNN spiking statistics for threshold initialization. However, the aforementioned methods were susceptible to information loss during SNN inference due to the use of hard reset spiking neurons, wherein the membrane potential is reset to zero in the event of a spike. han2020rmp demonstrated near lossless low-latency ANN-SNN conversion with soft reset neurons, wherein the potential is decreased by an amount equal to the neuronal spiking threshold. Subsequently, bu2023optimal showed that the threshold voltage could be learned, instead of configuring it heuristically post training, by replacing ReLU with the Quantization-Clip-Floor-Shift (QCFS) activation, which further minimizes the conversion error.
@@ 2.2 Temporally Efficient SNN Inference
In parallel with improving the ANN-SNN conversion process, there have been efforts to speed up SNN inference by optimizing the number of timesteps needed to achieve the best accuracy. Early-exit inference methods, which incorporate auxiliary classifiers at various stages in the model, have been shown to improve the inference efficiency of both ANNs and SNNs (panda2016conditional; huang2018multiscale; srinivasan2021bloctrain)li2023seenn proposed a unique early-exit approach for SNNs, wherein the number of inference timesteps is modulated based on the input complexity. In this work, we propose a novel methodology to achieve both precise and temporally efficient ANN-SNN conversion.
@@ 3 Methodology
3.1 Quantization-Clip-Floor-Shift (QCFS) Activation
The QCFS function, $\widehat{h}$, as proposed by bu2023optimal, can be formulated as
where $z^{l}$ is the input to the activation layer, $L$ is the quantization step, and $\lambda^{l}$ is the trainable parameter for the activation threshold. At the end of training, $\lambda^{l}=\theta_{n}$, where $\theta_{n}$ is the final threshold that is subsequently used for SNN inference. bu2023optimal showed that the expectation of conversion error reaches $\bm{0}$ when the shift amount, $\bm{\varphi}$, is $\bm{\frac{1}{2}}$.
@@ 3.2 Precise ANN-SNN Conversion (PASC)
Integrate-and-Fire (IF) layer. We consider the general case where the quantization step $L_{n-1}$ of the previous layer is different from that of the current layer ($L_{n}$). We initialize the layerwise threshold $\theta^{*}=\frac{\theta_{n}}{L_{n}}$, where $\theta_{n}$ is obtained from ANN training (refer to Section 3.1). The initial membrane potential is set to $\frac{\theta^{*}}{2}$, similar to prior works (bu2023optimal). We outline the procedure in Algorithm 1 for a generic layer, and Algorithm 2 for the input layer. We formally prove in Section 3.3 that the proposed algorithm enables precise ANN-SNN conversion. The algorithm for a generic layer is described below.
@@
    For the first $L_{n-1}$ timesteps, we perform the IF operation (detailed in Section 2.1) using the spike train obtained from the preceding layer. We use a spike counter to record the number of spikes produced at this stage.
@@
The output spike train generated by the PASC algorithm is effectively unrolled across $L_{n}$ timesteps by the subsequent layer. Therefore, we preserve the fact that after each layer, the number of inference timesteps is equal to the value of the quantization step, i.e. $T=L_{n}$. We note that Algorithm 2 is similar to, albeit simpler than, Algorithm 1, except that it receives the input image containing real-valued pixels (rather than binary spikes) over time. Feeding the input image to the SNN, without encoding the pixel intensities as spike trains, has been adopted in prior works to minimize accuracy loss during inference (li2023seenn; bu2023optimal).
@@
9:Set $mem(0)$ as the final membrane potential after the previous loop.
10:for $t\in\{1,2,\dots max(L_{n-1},L_{n})-1\}$ do
11:if $mem(t)\geq\theta^{*}$ then
12:  $s=s+\frac{\theta_{n}}{L_{n}}$ // Increment on Excitatory spike
13:  $mem(t)=mem(t)-\theta^{*}$
14:else if $mem(t)<0$ then
15:  $s=s-\frac{\theta_{n}}{L_{n}}$ // Decrement on Inhibitory spike
16:  $mem(t)=mem(t)+\theta^{*}$
17:Initialize a new SNN with threshold $\theta^{*}=\frac{\theta_{n}}{L_{n}}$, and initial membrane potential $mem(0)=s$, where $s$ is the spike count (scaled by $\theta^{*}$) computed previously.
18:Initialize an empty tensor $s^{l}$, with dimension $L_{n}\times N\times H\times W$, to generate the final spike train.
19:for $t\in\{1,2,\dots L_{n}\}$ do
[…]
22:return $s^{l}$, which will be the input for the subsequent MatMul layer.
@@
[…]
8:return $s^{l}$
MatMul layers. The MatMul layer performs either a convolution or a fully-connected matrix multiplication, followed by batch normalization. These operations are generally performed using Multiply-and-Accumulate (MAC) operations during ANN inference. It is important to note that the MatMul layers receive spike trains over $L_{n}$ timesteps from the preceding IF layer during SNN inference. The binary spike at each timestep is represented by $\{0,1\}\cdot\theta^{*}$, where the threshold $\theta^{*}=\frac{\theta_{n}}{L_{n}}$. We would like to point out that the quantity obtained by scaling a spike of unit magnitude with the threshold is referred to as postsynaptic potential in the literature (bu2023optimal). Therefore, all computations in MatMul layers are unrolled over time and performed on binary input. The former MAC operations are replaced by repeated additions carried out on the spike train.
@@ 3.4 AL: Adaptive Layerwise QCFS Activation
3:Compute the Skewness $g$ (Equation 4), the Van Der Eijk’s Agreement AA (refer Section 3.4.1) , and the kurtosis $K$ (Equation 5), using the histogram $\mathbf{H}$.
@@ 3.4.1 Van Der Eijk’s Agreement
We use Van Der Eijk’s agreement (vandereijk) to determine how many bins in the histogram $\mathbf{H}$ have a sufficient frequency. It is specified by
@@ 3.4.2 Skewness
Skewness is defined as the deviation of the given distribution from a normal distribution (skewnesskurtosis). It is given by
@@
We use the metric $g^{2}+1$, where $g$ is as defined in Equation 4, as adopted in prior works (sarlecoeff).
@@ 3.4.3 Kurtosis
where $k_{4}$ is the fourth moment, $k_{2}$ is the second moment, $n$ is the number of samples and $x_{i}$ are the samples whose mean is $\bar{x}$ (skewnesskurtosis). We interpret kurtosis in the following manner.
@@ 3.4.5 Adaptive Layerwise QCFS Activation based Training Methodology
We estimate the AL metric for all the QCFS layers of a pretrained ANN. We use 1-dimensional clustering (grønlund2018kmeans) to divide the layers into $\chi$ clusters based on the AL metric values, where $\chi$ is a user-defined parameter. The goal is to group QCFS layers with similar AL metric values and assign a suitable quantization step. We subsequently retrain the source ANN from the beginning. This approach is inspired by the Lottery Ticket Hypothesis (frankle2019lottery); the difference being that our subnetwork is not obtained by pruning, but rather by activation quantization. We employ a higher quantization step for the initial epochs to maximize the effect of the learning rate scheduler (loshchilov2016sgdr). We then configure the quantization step per layer and incrementally fine-tune the model until training converges, as outlined in Algorithm 4.